How many moles of H2 and HI will be present at equilbrium if 0.1570 mol HI are placed into a 1-L flask and allowed to react at a temperature where for the following reaction takes place:

2HI(g)  H2(g) + I2(g)K = 10.8



1
Expert's answer
2021-07-27T18:40:01-0400

Concentration of HI initially"=\frac{0.1570}{1}=0.1570M"


"2HI(g) = H_2(g) + I_2(g)" "K = 10.8"


"K=\frac{(H_2)(1_2)}{(HI)^2}" ="\frac{(x)(x)}{(0.1570-2x)^2}=10.8"


"K=[\frac{x}{0.1570+2x}]^2=10.8"


"K=[\frac{x}{0.1570-2x}]=3.29"


"0.1570-2x=3.29x"


"x= 0.03" hence the equlibrium


H2 =0.03